OpenStudy (anonymous):

Without any replacement, 2 cards are drawn from a well shuffled pack of 52 cards.What is the probability that neither a jack nor a card of spades is drawn?? Ans given [105/221]

OpenStudy (amistre64):

what is P(not a jack) ?

OpenStudy (anonymous):

1-4/52 = 48/52

OpenStudy (amistre64):

and what P(not a spade)?

OpenStudy (anonymous):

1-13/52 = 39/52

OpenStudy (amistre64):

now, if we add these, we have assumed they are independant, but they arent. some of one are some of the other and we are getting a double count what is the P(notjack and notspade?)

OpenStudy (anonymous):

there is one jack which belongs to the suite of spades, how to take that into account ?

OpenStudy (amistre64):

without replacement .. lets redo this ..

OpenStudy (amistre64):

the wording is confusing il admit

OpenStudy (amistre64):

Without replacement 2 cards are drawn What is the probability that neither a jack nor a card of spades is drawn j's' + s'j' + j'j' + s's' these are all the possible outcomes right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

if im wrong dont agree lol can you think of how this is bad?

OpenStudy (amistre64):

4 jacks, 13 spades, kick one thats already a jack 4+12 = 16 bad cards out of 52 favorable is therefore 36 out of 52 to pick 16 36 52 good it adds

OpenStudy (anonymous):

what i understand is we have two possibilities when we draw the jack first , we have 51 cards to choose the spade from , but what if the jack drawn was of spade ? second if we draw a spade first we have 51 cards to choose the jack from , but what is the spade drawn was a jack.

OpenStudy (amistre64):

P(s) * P(s|given first is an s) 36 35 -- * -- 52 51 does this give our answer?

OpenStudy (anonymous):

yes it does

OpenStudy (amistre64):

[105/221] its same

OpenStudy (amistre64):

so, how did we do it? you tell me

OpenStudy (anonymous):

could you explain what you did ?

OpenStudy (amistre64):

i explained it the best i could already, you tell me what your reading :)

OpenStudy (anonymous):

36 cards are those cards which are neither a jack nor a spade

OpenStudy (amistre64):

good, out of 52

OpenStudy (amistre64):

do you see how we get 36?

OpenStudy (anonymous):

yes , I got that, I did not understand 35/ 51 part

OpenStudy (amistre64):

when we pull a card, conditions change right?

OpenStudy (anonymous):

yes .. so we are pulling out two cards which are neither a jack nor a spade one after the other without replacement.

OpenStudy (amistre64):

given that we have removed one already from teh deck, and it was not a jack or spade ... how many good cards are left out of how many cards left overall?

OpenStudy (anonymous):

I get it now ... thank you !!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

i am closing the question

OpenStudy (amistre64):

thats fine, im looking for another one